OpenStudy (anonymous):

identifying - 2 x squared minus the quantity - 3 x squared plus 4x squared

OpenStudy (jhannybean):

Are you looking for how to solve this?

OpenStudy (jhannybean):

2x^2 - (3x^2 + 4x^2) Combine what is inside the parenthesis : 3x^2 + 4x^2 = (3+4)x^2 = 7x^2 Then subtract 2x^2 by 7x^2 2x^2 - 7x^2 (2-7)x^2 = ?

OpenStudy (jhannybean):

solve for that and you'll find your answer :)